# CrosGrpsABS: Cross-Attention over Syntactic and Semantic Graphs for Aspect-Based Sentiment Analysis in a Low-Resource Language

### v1

@@ III-H Classification Layer
IV Experiment and Result Analysis
IV-A Implementation Details
We conduct experiments on four Bangla ABSA benchmark datasets, Car (1,149 samples), Mobile Phone (975), Movie (800), and Restaurant (801), introduced by Hasan et al.[40], alongside English datasets from the SemEval 2014 Task 4 benchmark[1]. Table I and II report the data distributions for the Bangla and English datasets, respectively. The Bangla datasets, sourced from platforms like Facebook, YouTube, and online portals, include columns for Id, Comment, and Aspect category with Sentiment Polarity, while the SemEval English datasets cover the Laptop and Restaurant domains, providing sentences annotated with aspect terms and sentiment polarities to ensure fair comparison with state-of-the-art models. We perform cleaning steps to remove noise, correcting typographical errors, removing irrelevant symbols or URLs, and normalizing character encodings, ensuring a consistent and high-quality corpus. We then employ a pre-trained tokenizer to segment the text into subword units, leveraging its learned vocabulary to capture linguistic nuances and reduce out-of-vocabulary issues. Table III summarizes the hyperparameter settings and techniques: a batch size of 1, a maximum sequence length of $128$, $20$ training epochs at a learning rate of $2\times 10^{-5}$ for balanced convergence speed and stability, a single GATConv layer, and the AdamW optimizer with a weight decay of $1\times 10^{-2}$. To address overfitting, we apply early stopping callbacks and class weights to mitigate imbalance, and we set a hidden dimension of $768$ to capture linguistic representations. Finally, we assess performance using Test Accuracy (Acc) and Micro‑averaged F1 score (F1), providing a comprehensive evaluation of our aspect-based sentiment classification approach.
IV-B Findings
Table IV presents the comparative performance of our proposed CrosGrpsABS model using three different embedding techniques: BanglaBERT, mBERT [8], and RoBERTa [9], evaluated across four domain-specific Bengali ABSA datasets: Car, Mobile, Movie, and Restaurant. The results are reported in terms of accuracy (Acc) and micro F1 scores (F1). Among these embedding techniques, RoBERTa-based embeddings achieve the highest performance in most domains, notably in Car, Movie, and Restaurant datasets, with a peak accuracy of 88.74% and an F1 score of 88.69% on the Restaurant dataset. The mBERT embeddings yield competitive results, particularly in the Mobile domain, achieving the highest accuracy (86.49%) and micro F1 (86.32%). Conversely, the BanglaBERT embeddings lead to relatively lower performance, suggesting the advantage of utilizing multilingual transformer-based embeddings like RoBERTa and mBERT within the CrosGrpsABS framework.
IV-C Baseline Comparison
To evaluate the performance of our proposed CrosGrpsABS model, we conducted experiments on four domain-specific Bengali ABSA datasets: Car, Mobile, Movie, and Restaurant. These datasets are representative of real-world sentiment analysis tasks in Bengali. Table V presents a performance comparison between CrosGrpsABS and several baseline models, including transformer-based methods (RoBERTa-base, RoBERTa-large, RoBERTa, BanglaBERT, mBERT), graph-based models (GCN, Dual-GCN), and sequential neural networks (BiLSTM, BiGRU), as well as standard Transformer architecture. Our CrosGrpsABS model consistently outperforms all baseline models across all domains. Specifically, it achieves the highest accuracy and micro F1 score in the Restaurant dataset, with accuracy reaching 88.74% and a micro F1 of 88.69%. Among the baseline models, fine-tuned RoBERTa stands out as the strongest, consistently ranking second with a performance of 80.13% accuracy on Restaurant data. The zero-shot RoBERTa-base and RoBERTa-large models show lower performance across all domains, emphasizing the importance of task-specific fine-tuning. BanglaBERT, despite being language-specific, exhibits moderate results, demonstrating that pretraining alone may not suffice without further adaptation. Sequential models like BiLSTM and BiGRU, as well as graph-based models (GCN, Dual-GCN), show competitive but inferior performance, highlighting the effectiveness of our hybrid approach that integrates transformer-based embeddings with bidirectional cross-attention on syntactic and semantic graphs.
@@ IV-C Baseline Comparison
SemEval 2014 Task 4 results
Table VI provides a comprehensive comparison of our proposed CrosGrpsABS model against state-of-the-art baseline methods on the SemEval 2014 Task 4 datasets (Restaurant and Laptop). The baselines include recent advanced ABSA methods such as BERT, R-GAT+BERT, DGEDT+BERT, BART-GCN, T-GCN+BERT, DualGCN+BERT, SSEGCN+BERT, AG-VSR+BERT, TextGT+BERT, and DAGCN+BERT. In the Restaurant dataset, CrosGrpsABS achieves highly competitive results with an accuracy of 87.88% and micro F1 score of 87.93%, slightly below AG-VSR+BERT, which attains the highest accuracy at 89.96%. On the Laptop dataset, CrosGrpsABS clearly demonstrates superior performance by reaching the highest accuracy (86.43%) and micro F1 score (86.46%), surpassing all other baseline methods. These results validate the effectiveness and robustness of CrosGrpsABS across different datasets and domains, emphasizing its capability to compete with state-of-the-art ABSA approaches.
TABLE VIII: Examples from the Bengali ABSA Dataset with True Labels and Model Predictions.
IV-D Ablation Study
Table VII presents an ablation study conducted on our proposed CrosGrpsABS model, evaluating the contributions of different components across four Bengali ABSA datasets: Car, Mobile, Movie, and Restaurant. Specifically, we examine the effects of removing individual modules, including the syntactic graph, semantic graph, both graph branches, cross-attention, transformer embeddings, highway gating mechanism, aspect embeddings, and using a fixed adjacency matrix. The results demonstrate that the complete CrosGrpsABS model consistently outperforms all ablated variants, emphasizing the critical role of each component. Notably, removing the semantic graph significantly degrades performance, especially in the Mobile domain, dropping accuracy from 79.57% to 47.18%, which highlights the importance of semantic information. Excluding transformer embeddings also leads to notable performance reductions across all datasets, indicating transformers’ key role in encoding contextual information. Similarly, eliminating the syntactic graph results in decreased accuracy and F1 scores, particularly evident in the Car dataset, reducing accuracy from 70.89% to 58.70%. These findings collectively confirm the effectiveness and necessity of integrating both syntactic and semantic information with transformer embeddings and cross-attention mechanisms to achieve optimal performance.
IV-E Effect on GNN layers
@@
IV-F Visualization
@@
IV-G Case Study
Table VIII presents a detailed case study analysis, showcasing three challenging Bengali ABSA examples alongside predictions from Dual-GCN, RoBERTa, and our proposed CrosGrpsABS model. In the first example, both RoBERTa and CrosGrpsABS correctly identify negative sentiment for both Ambiance and Service, whereas Dual-GCN misclassifies Service as positive. The second example involves two positive aspects (Food and Price); Dual-GCN and RoBERTa each misjudge at least one of these aspects, while CrosGrpsABS accurately predicts both. Finally, the third example contains a contrasting statement (handbrake broken vs. normal brakes fine), and although Dual-GCN misclassifies the second Safety aspect, both RoBERTa and CrosGrpsABS correctly predict negative sentiment for both. These observations highlight CrosGrpsABS’s robustness in handling complex linguistic cues and mixed sentiment contexts within Bengali text.
V Conclusions
In this work, we proposed CrosGrpsABS, a novel hybrid model designed to enhance Aspect-Based Sentiment Analysis (ABSA) for low-resource Bengali. Our approach integrates transformer-based embeddings with bidirectional cross-attention on syntactic and semantic graph representations, combining local syntactic cues with global semantic contexts. Extensive experiments on domain-specific Bengali datasets and the SemEval 2014 benchmark demonstrate CrosGrpsABS’s superior performance compared to state-of-the-art baselines. Qualitative analyses, including attention visualizations and token-level importance scores, highlight the model’s interpretability. However, challenges remain, such as handling mixed sentiments, implicit emotional expressions, and reliance on accurate syntactic dependency parsing, which can be problematic for low-resource languages. Future improvements could include contrastive and context-aware learning for better handling of complex sentiments, multilingual transfer learning for broader generalization, and advanced graph-based techniques to refine syntactic and semantic dependencies. Expanding this methodology to other low-resource languages and domains could further enhance its practical applicability and advance ABSA research.

### v2

@@ 3.8 Classification Layer
4 Experimental Setup
4.1 Datasets
We conduct experiments on four public benchmark datasets for aspect-based sentiment analysis (ABSA) in Bangla, introduced by Hasan et al. [40], as well as on English datasets from the widely used SemEval 2014 Task 4 benchmark [1].
Our Bangla datasets consist of Car, Mobile Phone, Movie, and Restaurant, each provided in a dedicated Excel file (Car ABSA, Mobile phone ABSA, Movie ABSA, and Restaurant ABSA, respectively) with three columns: Id, Comment, and Aspect category, Sentiment Polarity. These datasets were compiled from platforms such as Facebook, YouTube, and various online portals. Specifically, Car has 1149 samples, Mobile Phone has 975, Movie contains 800, and Restaurant has 801, collectively offering a comprehensive resource for ABSA in Bangla. Table 1 presents our data distribution for the four domain-specific ABSA datasets in Bangla.
In addition, we evaluate on the SemEval 2014 Task 4 English datasets, which include the Laptop and Restaurant domains [1]. These datasets have been extensively used in previous ABSA research and serve as a standard benchmark for comparing model performance across languages. Each dataset contains sentences annotated with aspect terms and associated sentiment polarities (positive, negative, or neutral), ensuring a fair comparison with state-of-the-art ABSA approaches. Table 2 presents statistics of SemEval 2014 Task 4 datasets (Laptop and Restaurant) with sentiment distribution.
4.2 Data Preparation
In preparing the data, we first perform a series of cleaning steps to remove noise, such as correcting typographical errors, removing irrelevant symbols or URLs, and normalizing character encodings. This ensures a consistent and high-quality corpus for subsequent processing. Next, we employ a pre-trained tokenizer to segment the text into subword units, leveraging the tokenizer’s learned vocabulary to capture linguistic nuances and reduce out-of-vocabulary issues. By combining thorough data cleaning with an advanced, pre-trained tokenizer [9], we create a reliable foundation for downstream tasks while minimizing potential errors and preserving meaningful textual information.
4.3 Baselines
We compare our proposed CrosGrpsABS with several strong baseline models to evaluate its effectiveness. The baselines include Transformer-based models employing pre-trained embeddings: RoBERTa-base and RoBERTa-large evaluated under zero-shot conditions [9], fine-tuned RoBERTa [9], multilingual BERT (mBERT) [8], and BanglaBERT [41], a transformer explicitly pretrained on Bangla text. Additionally, we incorporate neural architectures such as Graph Convolutional Networks (GCN), Dual-GCN, Bidirectional LSTM (BiLSTM), and Bidirectional GRU (BiGRU), all enhanced with Transformer-based embeddings. These models provide comprehensive comparative benchmarks, reflecting state-of-the-art approaches commonly employed in ABSA tasks.
Evaluation on SemEval 2014 Task 4
We further evaluate the performance of our proposed model, CrosGrpsABS, on the widely-used SemEval 2014 Task 4 benchmark datasets, specifically the Laptop and Restaurant domains. To provide a rigorous comparison, we include several state-of-the-art baselines previously established in ABSA literature, including BERT [8], R-GAT+BERT [34], DGEDT+BERT [35], BART-GCN [36], T-GCN+BERT [18], DualGCN+BERT [19], SSEG-CN+BERT [37], AG-VSR+BERT [18], TextGT+-BERT [18], and DAGCN+BERT [38]. These strong baselines enable a comprehensive assessment of our model’s capability to generalize across different languages and datasets.
@@ 4.3 Baselines
4.4 Ablation Study
To systematically evaluate the contribution of each component within our proposed model, CrosGrpsABS, we conduct extensive ablation experiments under various settings. Specifically, we investigate the impact of removing key components, including the syntactic graph (No Syntax Graph), semantic graph (No Semantic Graph), both graph branches entirely (No Graph Branches), the cross-attention mechanism (No Cross-Attention), Transformer embeddings (No Transformer), the highway gating mechanism (No Highway Gate), aspect embeddings (No Aspect Embedding), and utilizing a fixed adjacency matrix (Fixed Adjacency). These analyses provide deeper insights into the effectiveness of individual modules and their interactions within our proposed architecture.
4.5 Computational Resources
All experiments were conducted using Python 3.12.9 and PyTorch 2.4, utilizing an NVIDIA GeForce RTX 2060 GPU with 16 GB of RAM. This hardware and software configuration ensured efficient training and evaluation of our models, allowing reproducible and reliable comparisons across various baseline methods and proposed architectures.
4.6 Evaluation Metrics
To assess the performance of our proposed model and baseline methods, we utilize two widely adopted evaluation metrics: Test Accuracy (Acc) and Micro-averaged F1 score (F1). Accuracy provides a straightforward measure of overall correctness, while the Micro F1 score offers a balanced evaluation by considering both precision and recall across all classes, particularly beneficial for datasets with class imbalance. These metrics collectively enable comprehensive and reliable evaluation of our aspect-based sentiment classification approach.
4.7 Hyperparameter Settings
Table 3 summarizes the hyperparameter settings and additional techniques used in our experiments. We evaluated multiple batch sizes and sequence lengths, identifying a batch size of 1 and a maximum sequence length of 128 as optimal. Training for 20 epochs at a learning rate of $2\times 10^{-5}$ balanced convergence speed and stability. We chose a single GATConv layer for simplicity and efficiency, while AdamW (with a weight decay of $1\times 10^{-2}$) served as our optimizer. To address potential overfitting, we used early stopping callbacks, and we applied class weights to mitigate imbalance. Finally, a hidden dimension of 768 provided ample capacity for capturing nuanced linguistic representations.
5 Results and Discussions
Table 4 presents the comparative performance of our proposed CrosGrpsABS model using three different embedding techniques: BanglaBERT, mBERT [8], and RoBERTa [9], evaluated across four domain-specific Bengali ABSA datasets: Car, Mobile, Movie, and Restaurant. The results are reported in terms of accuracy (Acc) and micro F1 scores (F1). Among these embedding techniques, RoBERTa-based embeddings achieve the highest performance in most domains, notably in Car, Movie, and Restaurant datasets, with a peak accuracy of 88.74% and an F1 score of 88.69% on the Restaurant dataset. The mBERT embeddings yield competitive results, particularly in the Mobile domain, achieving the highest accuracy (86.49%) and micro F1 (86.32%). Conversely, the BanglaBERT embeddings lead to relatively lower performance, suggesting the advantage of utilizing multilingual transformer-based embeddings like RoBERTa and mBERT within the CrosGrpsABS framework.
Figure 4: Attention heatmap illustrating the token-to-token attention weights for a sample Bengali sentence. Higher-intensity regions (yellow) indicate stronger attention signals between corresponding tokens.
5.1 Baseline Comparison
Table 5 presents the comparative performance of our proposed CrosGrpsABS model against multiple baseline models on four domain-specific Bengali ABSA datasets: Car, Mobile, Movie, and Restaurant. The baseline methods include transformer-based embeddings (RoBERTa-base and RoBERTa-large evaluated under zero-shot settings, RoBERTa, BanglaBERT [41], mBERT [8, 9]), as well as graph-based models (GCN, Dual-GCN), sequential neural networks (BiLSTM, BiGRU), and standard Transformers architecture. Our CrosGrpsABS model consistently achieves the highest performance across all domains, significantly outperforming all baseline models. Specifically, CrosGrpsABS attains its best results in the Restaurant dataset, achieving an accuracy of 88.74% and a micro F1 of 88.69%. Among the baselines, fine-tuned RoBERTa shows robust performance and emerges as the strongest baseline, consistently ranking second-best, notably with 80.13% accuracy on Restaurant data. The lower performance of zero-shot RoBERTa-base and RoBERTa-large across all domains highlights the importance of task-specific fine-tuning. Meanwhile, BanglaBERT exhibits moderate performance, emphasizing that even language-specific pretraining alone may not suffice without further adaptations. Sequential models such as BiLSTM and BiGRU, and the graph-based models (GCN, Dual-GCN) demonstrate competitive but inferior performance, validating the effectiveness of our hybrid approach that integrates transformer-based embeddings with bidirectional cross-attention on syntactic and semantic graphs.
@@ 5.1 Baseline Comparison
SemEval 2014 Task 4 results:
Table 6 provides a comprehensive comparison of our proposed CrosGrpsABS model against state-of-the-art baseline methods on the SemEval 2014 Task 4 datasets (Restaurant and Laptop). The baselines include recent advanced ABSA methods such as BERT, R-GAT+BERT, DGEDT+BERT, BART-GCN, T-GCN+BERT, DualGCN+BERT, SSEG-CN+BERT, AG-VSR+BERT, Tex-tGT+BERT, and DA-GCN+BERT. In the Restaurant dataset, CrosGrpsABS achieves highly competitive results with an accuracy of 87.88% and micro F1 score of 87.93%, slightly below AG-VSR+BERT, which attains the highest accuracy at 89.96%. On the Laptop dataset, CrosGrpsABS clearly demonstrates superior performance by reaching the highest accuracy (86.43%) and micro F1 score (86.46%), surpassing all other baseline methods. These results validate the effectiveness and robustness of CrosGrpsABS across different datasets and domains, emphasizing its capability to compete with state-of-the-art ABSA approaches.
Table 8: Examples from the Bengali ABSA Dataset with True Labels and Model Predictions.
5.2 Ablation Study
Table 7 presents an ablation study conducted on our proposed CrosGrpsABS model, evaluating the contributions of different components across four Bengali ABSA datasets: Car, Mobile, Movie, and Restaurant. Specifically, we examine the effects of removing individual modules, including the syntactic graph, semantic graph, both graph branches, cross-attention, transformer embeddings, highway gating mechanism, aspect embeddings, and using a fixed adjacency matrix. The results demonstrate that the complete CrosGrpsABS model consistently outperforms all ablated variants, emphasizing the critical role of each component. Notably, removing the semantic graph significantly degrades performance, especially in the Mobile domain, dropping accuracy from 79.57% to 47.18%, which highlights the importance of semantic information. Excluding transformer embeddings also leads to notable performance reductions across all datasets, indicating transformers’ key role in encoding contextual information. Similarly, eliminating the syntactic graph results in decreased accuracy and F1 scores, particularly evident in the Car dataset, reducing accuracy from 70.89% to 58.70%. These findings collectively confirm the effectiveness and necessity of integrating both syntactic and semantic information with transformer embeddings and cross-attention mechanisms to achieve optimal performance.
5.3 Effect on GNN layers
@@
5.4 Visualization
Figure 4 illustrates an attention heatmap generated by our CrosGrpsABS model, demonstrating token-to-token attention weights for a sample Bengali sentence. The heatmap visualizes attention intensity between pairs of tokens, with higher-intensity regions (indicated by yellow) representing stronger attention signals. This visualization highlights how CrosGrpsABS effectively captures the relationships between different words within the sentence, providing insights into how syntactic and semantic information is integrated through cross-attention mechanisms. Such attention visualization further aids interpretability by allowing analysis of the model’s reasoning process when identifying sentiment-bearing aspects in Bengali texts.
@@ 5.4 Visualization
5.5 Case Study
Table 8 presents a detailed case study analysis, showcasing three challenging Bengali ABSA examples alongside predictions from Dual-GCN, RoBERTa, and our proposed CrosGrpsABS model. In the first example, both RoBERTa and CrosGrpsABS correctly identify negative sentiment for both Ambiance and Service, whereas Dual-GCN misclassifies Service as positive. The second example involves two positive aspects (Food and Price); Dual-GCN and RoBERTa each misjudge at least one of these aspects, while CrosGrpsABS accurately predicts both. Finally, the third example contains a contrasting statement (handbrake broken vs. normal brakes fine), and although Dual-GCN misclassifies the second Safety aspect, both RoBERTa and CrosGrpsABS correctly predict negative sentiment for both. These observations highlight CrosGrpsABS’s robustness in handling complex linguistic cues and mixed sentiment contexts within Bengali text.
6 Conclusion
In this work, we proposed CrosGrpsABS, a novel hybrid model designed specifically to enhance Aspect-Based Sentiment Analysis (ABSA) for the low-resource Bengali language. Our approach uniquely integrates transformer-based embeddings with bidirectional cross-attention mechanisms on syntactic and semantic graph representations, effectively combining local syntactic cues with global semantic contexts. Extensive experiments conducted across multiple domain-specific Bengali datasets, as well as the SemEval 2014 benchmark datasets, have demonstrated the superior performance of CrosGrpsABS compared to several state-of-the-art baselines. Moreover, qualitative analyses, such as attention visualizations and token-level importance scores, have provided valuable insights into the interpretability and reasoning mechanisms of our proposed framework.
However, despite the overall effectiveness of CrosGrpsABS, certain limitations remain evident. Specifically, the model sometimes struggles with sentences expressing mixed sentiments, implicit emotional expressions, or contexts involving contrasting or comparative aspects. Such cases often require deeper semantic comprehension and nuanced reasoning beyond the model’s current capabilities. Another limitation pertains to the model’s reliance on accurate syntactic dependency parsing, which can be challenging for languages with limited linguistic resources, potentially affecting the model’s performance in less structured scenarios. Potential improvements include developing advanced contrastive and context-aware learning techniques to better handle sentences with implicit or mixed sentiments. Additionally, exploring multilingual or cross-lingual transfer learning approaches could further improve the robustness and generalizability of ABSA methods for low-resource languages. Future research could also investigate more sophisticated graph-based techniques, such as dynamic adjacency mechanisms or adaptive graph construction, to further refine the modeling of syntactic and semantic dependencies. Finally, expanding this methodology to other low-resource languages and diverse domains could significantly enhance its practical applicability, thereby advancing the broader field of ABSA.
Conflits of Interest
The authors declare that they have no conflict of interest.
[…]